OpenStudy (anonymous):

Find all solutions of the equation in the interval [0,2pi] A) 2sin3x-1=0

OpenStudy (cwrw238):

sin 3x = 0.5 3x = pi/3 in first quadrant can you continue?

OpenStudy (anonymous):

so pi/6

OpenStudy (cwrw238):

no pi/3 / 3 = pi/9 thats one solution

OpenStudy (anonymous):

NO ITS PI OVER 6

OpenStudy (cwrw238):

as we want x in the interval [0,2pi] we have to consider 3x between 0 and 6pi.

OpenStudy (anonymous):

can u draw this out pleaz

OpenStudy (anonymous):